• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 00:03:35 on August25,2006

    Entry number 182346

    keyword=shift summary

    beam back at 4:20 ; 20 mkA on C-12 target for kin #6
    start run 3259/21538 at 16:28 to repeat C-12 calibration
    due to small amplitudes in some S1 and S0- start HV adjustment
    3264 is first run to use - S0 voltage 1850 V
    run 3264/21543 - C-12 production 1M
    run 3265/21544 - C-12 production 1M
    run 3266/21545 - C-12 production 1M - raster ON
    run 3267/21546 - C-12 production 1M - raster OFF
    run 3268/21547 - 4cm Dummy 10 mkA - raster ON